# Learning to Visually Connect Actions and their Effects

### v1

@@ 3.2 Baseline models
This framework consists of a state change computer (SCC) head, and an action encoder head (Figure 2(e)). SCC head takes in the scene information abstracted by backbone from initial and final states; and analyzes the two sets of features and computes the state-change occurring between the two states. To do this, SCC head concatenates the two set of features and processes them through a multi-layer perceptron, yielding a compact feature vector representing the state change. Operations like addition and subtraction can be alternatives to our deep state-change computer. However, we hypothesize that such operations are more likely to work where camera motion is absent or minimal. In the presence of camera motion, the natural pixel-to-pixel correspondence no longer holds, and such hard-coded operations would underperform, while more flexible solutions like our deep SCC would be better able to handle such noisy cases. The Action encoder extracts information from action clip. During training, we bring close together the state change and the correct action representations, while pushing apart state change and incorrect action representations. We found alternating between optimizing state and action heads to improve the performance. We hypothesize that this process can be viewed as the SCC learning to imagine action from states and framework making an analogy between imagined action and actual action options. Thus, we term this model as analogical reasoning model.
4 Task Formulation-II: Self-Supervised Action Selection
Figure 3: Self-supervised pretext task based on connecting actions and effects. Please zoom-in to view better. Applying action can bring the scene from its initial state to the final state, while applying temporally reversed version of it cannot. State-encoders are frozen; state-encoders and discriminator can be discarded after training—only the action encoders are retained for further usage.
@@ 4 Task Formulation-II: Self-Supervised Action Selection
The concept of connecting actions and their effects is closely related to the idea of the sensorimotor loop and the reciprocal relationship between vision and movement/action in human development [8, 7]. This reciprocal relationship involves learning via self-supervision. Similarly, we hypothesize that useful action representations can be learnt by deep video understanding models via learning to visually connect actions and their effects in a self-supervised way from unlabeled videos.
Objective.
@@
Given an unlabeled video of some human action, first and last frames are chosen to be representatives of initial and the final states of a scene. The remaining middle portion of the video is considered as the action(-clip). Now, we design a self-supervised task based on the observation that while applying action to the initial state would result in the final state; applying temporally-reversed version of the same action to the initial state would not result in the given final state. Concept along with implementation framework is shown in Figure 3. Note that the framework is a slightly modified version of Analogical Reasoning model from § 3 and explained further below.
Framework.
@@
First, let us introduce two ordered paired representations: 1) forward state-change: {initial state representation, final state representation}—order-preserving concatenation of the state representations; 2) backward state-change: {final state representation, initial state representation}. In order to avoid feature collapse, state representations are extracted using a SSL-pretrained feature extractor [49]. Note that, using SSL-pretrained feature extractor for downstream SSL methods is an accepted common practice as done in, e.g., [45, 37, 12]. During training the action network is tasked with making forward-action representation as similar to the forward state-change, and backward-action representation similar to the backward state-change. Once the action model is trained, it is used for downstream task such as action recognition. Discriminator is discarded after training.
5 Task Formulation-III: Self-Supervised Action Specification
@@
Figure 4: Self-supervised action specification. Zoom-in to view better. MDR: more directly related effect; LDR: less directly related effect. Degree of how directly related an effect is given on a scale of 1.0-0.0 written above effects. Farther the effect-frame from the action, lesser directly related it is to the action. Here we have shown cropped video frames to focus on the diver—in practice, we use the entire frame, and the network learns to focus on the diver through Action Specification-based self-supervision.
Objective.
@@
Action Specification aims to achieve a more nuanced understanding by focusing on discerning subtle effects in close proximity and determining the “distance” of an effect-frame from the action applied to an initial state. This task is formulated based on the observation that, when a short action is applied to an initial state, the frame immediately following the completion of the short action represents the most direct effect. As the video progresses, subsequent frames gradually exhibit a diminishing connection to the effect of the initial short action. Note that, the distance of an effect frame being a measurable and observable aspect, is used as a proxy to reflect the strength of the connection that effect-frame to the action.
@@
During training, the network learns to determine the “distance” (or equivalently, degree of relatedness, DR) of effect-frames from actions. Specifically, each training sample consists of an initial state, a short action clip, and a set of $K$ temporally ordered frames following the action, which form effect-frames. Each of the $K$ effect-frames are assigned a proxy groundtruth label which indicates the distance of an effect-frame from the end of the action clip. Specifically, in a series of $K$ temporally-ordered frames, $k$-th frame is assigned a value of $1-k/(K-1)$, where $k$ ranges from 0 to $K-1$. This way, the closest or the most directly-related frame is assigned a proxy-label of 1, while the farthest effect-frame is assigned a proxy-label of 0. The task for the network is to predict these proxy-labels as accurately as possible. Framework is shown in Figure 4. Network parameters are optimized through a regression loss (L2) function. Note that groundtruth labels ($01:1/K:$ 0) are artificially generated and are arbitrary—any other scale can be used as long as it can express how far is an effect-frame from action clip. We use unlabeled videos and artificially generated groundtruth to learn representations using this self-supervised pretraining task.
@@
In the process of solving this task, we hypothesize that the model learns to focus on intricate details, such as identifying the specific maneuvers executed by the athlete and understanding the consequent sequence of poses. Additionally, it learns to discern how the athlete’s position evolves as a direct outcome of these maneuvers. These details have an overlap with the elements of interest in the task of action quality assessment (AQA). Therefore, we hypothesize that representations learnt in solving our SSL CATE Action Specification should help with downstream task of AQA. To validate this, the self-supervised model is tested on action quality score prediction task, where the model scores how well an action was performed. A brief introduction to AQA task and implementation details are provided in the supplementary materials.
@@ 6 Experiments
In this section, we conduct the following experiments:
  • Benchmarking on Action Selection task
  • Probing into where the models need to look or pay attention in order to solve Action Selection task
  • Leveraging Action Selection task for improving on action anticipation task
  • Self-supervised representation learning by learning to connect actions and effects
  • Self-supervised action specification
  • Pose-encoding representations from solving self-supervised action specification
6.1 Cross Sample Action Selection
Model Performance
@@ 6.1 Cross Sample Action Selection
AEXformer 54.65
Explicit Disentangler 52.25
@@
Table 1: Performance of various approaches at Action Selection task.
Implementation details.
@@
All the models are implemented using PyTorch [32], and optimized using ADAM optimizer [21] with a learning rate of 1e-4. We train all the models using CATE train set, and select the best performing iteration based on the validation set. The performance of the best model is then verified on the test set. Noting the state-of-the-art performance of Video Swin model [25] on various video understanding tasks, we use a Kinetics pretrained version of it as the backbone feature extractor. Note that CATE task and methods are not limited to video swin model, and future work may use other models.
Answer evaluation.
@@
For the model’s answer to be considered correct it has to, from the given options select only one option that matches the ground truth based on the highest similarity. In the scenario where all options return equal similarities, it it deemed to have failed to answer correctly. We use accuracy to evaluate the performance metric.
@@
In this experiment, we measure the performance of various baseline approaches on the CATE task using the full CATE dataset. The results are summarized in Table 1. The Naive approach performs much better than random weights/chance accuracy. The approach treating actions as a transformation (AT) performs on par with the naive model. However, transformation-based approaches could be made more effective by simplifying the learning process and utilizing element-wise learnable weights for the interactions between the initial state and action vectors (MoRISA). We observe further improvements by learning a linear space for state-action interactions (AEXformer). Although it is a simple approach, it is found to be effective for connecting actions and effects. Learning to explicitly tease out or disentangle the ’transformation’ from actions and states works better than naive and AT approaches; however, it performs slightly inferior to AEXformer. While both approaches can achieve disentanglement in some form, empirically we observe that implicitly disentangling in the process of learning a linear space might be a better alternative than explicit disentanglement via feature partition and swapping. The Analogical Reasoning model performed the best. However, it is worth noting that training it is a bit more complicated than other models, as we had to resort to alternating between optimizing state change encoders and action encoders. Nonetheless, we observed that it achieves peak performance faster than other models. AEXformer and Explicit Disentangler take the longest to achieve peak performance.
Human baseline.
To put the performance of our baseline models into perspective, we conduct a study to assess human performance. We prepare a smaller subset of 100 questions and observe the performance of three undergraduate students on it. The average human performance is then compared with our model’s performance on this subset—see Table 1. We observe that humans performed significantly better than all of our baseline models on this seemingly easy task of connecting actions and effects.
Qualitative results.
We have visualized where our models “look” or focus when solving our visual action-effect task in Figure 5.
@@
Naive model Ours (Cross Sample)
@@
Figure 5: Probing where the model pays attention when connecting action and its effects.. (a) Action is ‘putting something close to something’. (b) Where our model attends vs where action recognition model attends. Our model focuses on state changes and driving action and effects; notice the avocado and coin being tracked. Action recognition without reasoning module seems to be doing simple matching based on texture. Please zoom in to view better.
Model Performance
Zatsarynna et al. [48] 13.39
Zatsarynna et al. [48] w Goal Consistency 13.93
Ours 18.80
Ours w generated effect 19.02
Table 2: Utility of generating effect in action anticipation task on COIN dataset.
Action Anticipation.
We explore the utility of effect generation in action anticipation. Following the literature [48], we conduct an action anticipation experiment on the COIN dataset. The task is to predict the next action one second into the future. Drawing inspiration from [14], we hypothesize that being able to generate unseen futures can aid by providing missing information. Thus, we generate this missing information using our AEXformer, which is already trained on our CATE Action Selection dataset and then frozen. At this stage, we simply use it as a “future feature extractor.” We then integrate this future feature with the current action feature and learn a simple decoder to predict the next action one second into the future. We also consider a case where we do not leverage the future feature. Frameworks for both approaches are shown in Figure 6. The top-5 mean recall results are presented in Table 2. It is evident that the future feature does help with the action anticipation task as well.
Figure 6: Action anticipation framework. (a) action anticipation model uses the “future feature” generated by our AEXformer; (b) model does not use the future feature.
6.2 Leveraging CATE task to learn visual representations from unlabeled videos.
After pretraining with the self-supervised action selection task (section 4) on the UCF101 [35] train set without using any labels, we freeze the backbone (C3D architecture [39]). No fine-tuning is performed. The frozen backbone is then used as a feature extractor. Following a standard practice in the literature [16, 40], we conduct a nearest neighbor retrieval experiment. Quantitative results for the nearest neighbor retrieval experiment are presented in Table 3. Our CATE model is compared with other state-of-the-art SSL video representation learning approaches. We observe that SSL-CATE performs the best overall in terms of Top-1 accuracy, ranking second best in the remaining metrics, only outperformed by STS. These results highlight the significance of CATE in effectively supporting the learning of general action representations.
@@
Table 3: Nearest Neighbor retrieval quantitative results on UCF101. Best performing method is bold-faced, and the second best method is underlined.
We further probe into what the model might be learning by retrieving videos similar to the randomly chosen query videos. Results are presented in Figure 7. We observe that the retrieved results are mostly from the correct action class, suggesting that CATE encourages the model to learn the tracking of important human body parts and objects. Upon further investigation of some cases where the retrieval is from the incorrect action class, we found that motion patterns in the retrieved videos can potentially result in the same state change as the action in the query videos. For example, in archery and playing cello, pulling the arrow in archery and moving the bow/stick on the cello involve similar hand movements from humans. We further compare the results with a fully-supervised model. Qualitatively, we found that the CATE-pretrained retrieved results are on par with a model trained with full action class label supervision.

### v3

@@ 3.2 Cross-Sample Analogical Reasoning Model for Action Selection
Framework. Actions are represented as video clips, while states by very short clips. Unless specified otherwise, we use a backbone model ($\mathcal{E}$) pretrained on a large-scale action recognition dataset to extract features for actions and states. Subsequently, various modules mentioned in the following operate on these features. The framework consists of a state change computer (SCC) head, and an action encoder head. SCC head takes in the scene information abstracted by backbone from initial and final states; and analyzes the two sets of features and computes the state-change occurring between the two states. To do this, the SCC head concatenates the two sets of features and processes them through a multi-layer perceptron, yielding a compact feature vector representing the state change. Operations like addition and subtraction can be alternatives to our deep SCC. However, we hypothesize that such operations are more likely to work where camera motion is absent or minimal. In the presence of camera motion, the natural pixel-to-pixel correspondence no longer holds, and such hard-coded operations would underperform, while more flexible solutions like our deep SCC would be better able to handle such noisy cases. The Action encoder extracts information from action clip. We hypothesize that the SCC focuses on object placement and pose; while the Action Encoder focuses on motion and tracks objects as shown in Figure 5.
Training. We use counterfactual reasoning [4], maximizing similarity between state changes & correct actions while minimizing it for incorrect actions. This approach outperformed only maximizing similarity with correct actions without discriminating incorrect actions in our experiments.
Optimization strategy. We found that alternating between optimizing state- & action-heads improves performance.
We hypothesize that this framework involves the SCC learning to imagine actions from states and making analogies between imagined and actual actions. Thus, we term this model the Analogical Reasoning model.
4 Emergent Representations from using
Action Selection as Self-Supervision
@@ 4 Emergent Representations from using
The concept of connecting actions and their effects is closely related to the idea of the sensorimotor loop and the reciprocal relationship between vision and movement/action in human development [10, 9]. This reciprocal relationship involves learning via self-supervision. Similarly, we hypothesize that useful action representations can be learnt by deep video understanding models via learning to visually connect actions and their effects in a self-supervised way from unlabeled videos. Towards that end, in this section, we show how CATE can be a useful source of intrinsic self-supervision for visual feature learning.
Figure 3: Self-supervised pretext task based on connecting actions and effects. Please zoom-in to view better. Applying action can bring the scene from its initial state to the final state, while applying a temporally reversed version of it cannot. State-encoders are frozen; state-encoders and discriminator can be discarded after training—only the action encoders are retained for further usage.
@@
Self-Supervised Task Design. Given an unlabeled video of some human action, the first and last frames are chosen to be representatives of the initial and the final states of a scene. The remaining middle portion of the video is considered as the action(-clip). Now, we design a self-supervised task based on the observation that applying action to the initial state would yield the final state; however, applying a temporally-reversed version of the same action to the initial state would not yield the given final state. In other words, the video understanding model is tasked with selecting the action which can take the scene from its initial state to final state—forward version of the action would, while temporally-reversed would not. We hypothesize in the process of solving this task, the model can learn useful video representations. This task has some resemblance with works like [30, 7], however, these works randomly shuffle frames, while our task operates specifically on the {initial state, action, final state} triplet. With random shuffling, potentially low-level cues could be leveraged; while our formulation offers connecting and discerning actions and effects at a semantically higher level. The concept along with implementation framework is shown in Figure 3. Note that the framework is a slightly modified version of Analogical Reasoning model from § 3 and explained further below.
@@
Framework. First, let us introduce two ordered paired representations: 1) forward state-change: {initial state representation, final state representation}—order-preserving concatenation of the state representations; 2) backward state-change: {final state representation, initial state representation}. To avoid feature collapse, state representations are extracted using a SSL-pretrained feature extractor [56]. Note that, using SSL-pretrained feature extractor for downstream SSL methods is an accepted common practice as done in, e.g., [53, 44, 14]. During training, the Action Encoder learns to makes forward-action representation similar to forward state-change, and backward-action representation similar to backward state-change. Once trained, the Action Encoder is used for downstream tasks like action recognition, and the discriminator is discarded.
5 Problem of Effect-Affinity Assessment
@@
Figure 4: Self-supervised Effect-Affinity Assessment. Zoom-in to view better. MDR: more directly related effect; LDR: less directly related effect. Degree of how directly related an effect is given on a scale of 1.0 to 0.0 (written above Effect-frames). Farther the effect-frame from the action, lesser directly related it is to the action. Here we have shown cropped video frames to focus on the divers’ pose; in practice, we use the entire frame, & the network learns to focus on the diver through Effect-Affinity Assessment-based self-supervision.
@@
Effect-Affinity Assessment aims to achieve a more nuanced understanding by focusing on discerning subtle effects in close proximity and determining the “distance” of an effect-frame from the action applied to an initial state. This task is formulated based on the observation that, when a short action is applied to an initial state, the frame immediately following the completion of the short action represents the most direct effect. As the video progresses, subsequent frames gradually exhibit a diminishing connection to the effect of the initial short action. Note that, the distance of an effect frame being a measurable and observable aspect, is used as a proxy to reflect the strength of the connection that effect-frame to the action.
@@
Framework. During training, the network learns to determine the “distance” (or equivalently, degree of relatedness, DR) of effect-frames from actions. Specifically, each training sample consists of an initial state, a short action clip, and a set of $K$ temporally ordered frames following the action, which form effect-frames. Each of the $K$ effect-frames are assigned a proxy groundtruth label which indicates the distance of an effect-frame from the end of the action clip. Specifically, in a series of $K$ temporally-ordered frames, $k$-th frame is assigned a value of $1-k/(K-1)$, where $k$ ranges from 0 to $K-1$. This way, the closest or the most directly-related frame is assigned a proxy-label of 1, while the farthest effect-frame is assigned a proxy-label of 0. The task for the network is to predict these proxy-labels as accurately as possible. Framework is shown in Figure 4. Network parameters are optimized through a regression loss (L2) function. Note that ground-truth labels ($01:1/K:$ 0) are artificially generated and are arbitrary—any other scale can be used as long as it can express how far is an effect-frame from the action clip. We use unlabeled videos and artificially generated ground-truth to learn representations using this self-supervised pretraining task.
@@
In the process of solving this task, we hypothesize that the model learns to focus on intricate details, such as identifying the specific maneuvers executed by the athlete and understanding the consequent sequence of poses. It also learns to discern how the athlete’s position evolves as a direct outcome of these maneuvers. These details have an overlap with the elements of interest in the task of action quality assessment (AQA). Thus, we hypothesize that representations learnt in solving our SSL CATE Effect-Affinity Assessment should help with downstream task of AQA. To validate this, the self-supervised model is tested on action quality score prediction task, where the model scores how well an action was performed. A brief introduction to AQA task and implementation details are provided in the Appendix.
@@ 6 Experiments
6.1 Cross-Sample Action Selection
We start by evaluating various baseline approaches on the Action Selection task/dataset.
Baselines. As the proposed task is new with no established methods/benchmarks, we design & explore various baseline methods to connect actions & effects. Below is a summary of the models; further details in the Appendix. Dataset & Codes for all models will be publicly released.
  • Naive baseline: cosine similarity between the average of the initial and final state features and each of the action options is computed. Action option with the highest similarity score is selected as the correct answer. This model involves no training, pretrained action recognition model is used as the state and action feature extractor.
  • VideoMAE[46, 6]: a model that reconstructs missing parts of video frames to learn video representations.
  • CLIP[38]: model learns to understand images and text together by training on a wide variety of internet data to create strong, versatile visual and textual representations.
  • VideoChat2[24]: strong multimodal foundational model (CVPR 2024) that is simultaneously trained on 20 challenging video understanding tasks. We measure its zeroshot performance.
  • Actions as Transformations: inspired by [50], we develop a model where a dense network computes a transformation vector from an action clip. The Hadamard product of the initial state vector & the transformation vector yields the final state vector. Inference: action option producing the highest similarity with the final state feature is chosen as the correct answer. Inference remains same for all baselines.
  • MoRISA: we leverage a bilinear model [26] to capture rich interactions between the initial state and the action to generate the final state in feature space.
  • LinSAES: we learn a linear, disentangled state-action embedding space, where adding action vector to the initial state vector, transitions it to final state vector.
  • Connecting via Swapping: We develop an encoder-decoder model. Encoders distill State-transformation ‘codes’ (STC) in two ways: 1) from pairs of initial and final states, and 2) from action clips. The decoder learns to construct the final state from the initial state and the STC. By training the decoder to recreate the final state using STCs from both state pairs and action clips, the model learns to understand how actions cause changes and can predict the outcomes of actions.
Model Accuracy(%)
@@ 6.1 Cross-Sample Action Selection
LinSAES 54.65
Connecting via Swapping 52.25
@@
Table 1: Performance of models on Action Selection task.
@@
Implementation details. All the models are implemented using PyTorch [37], and optimized using ADAM optimizer [23] with a learning rate of 1e-4. We train all the models using CATE train set and select the best-performing iteration based on the validation set. The performance of the best model is then verified on the test set. Noting the state-of-the-art performance of Video Swin model [28] on various video understanding tasks, we use a Kinetics pretrained version of it as the backbone feature extractor. Note that CATE task and methods are not limited to video swin model, and future work may use other models.
@@
Answer evaluation. For the model’s answer to be considered correct it has to, from the given options select only one option that matches the ground truth based on the highest similarity. In the scenario where all options return equal similarities, it is deemed to have failed to answer correctly. We use accuracy to evaluate the performance metric.
@@
Quantitative results. The results are summarized in Table 1. The Naive approach performs much better than random weights/chance accuracy. VideoMAE performs similar to Naive approach, suggesting that Action Selection and VideoMAE are different. Interestingly, foundational models CLIP and VideoChat2 performed inferior to naive model. This maybe because while MLLM/VLM have achieved strong results, they maybe less sensitive to finegrained details like object placement, pose, motion, which are crucial in our problem. This also suggests that CATE is a fundamentally different task than existing video understanding tasks, and foundational models can benefit from including it in their training suite. The Actions as Transformations (AT) approach performs on par with the naive model. However, transformation-based approaches could be made more effective by simplifying the learning process and utilizing per-element learnable weights for the interactions between the initial state and action vectors (MoRISA). We observe further improvements by learning a linear space for state-action interactions (LinSAES). Although it is a simple approach, it is found to be effective for connecting actions and effects. Connecting via Swapping which involves learning to explicitly tease out or disentangle the ‘transformation’ from actions and states works better than naive and AT approaches; however, it performs slightly inferior to LinSAES. While both approaches can achieve disentanglement in some form, empirically we observe that implicitly disentangling in the process of learning a linear space might be a better alternative than explicit disentanglement via feature partition and swapping. However, the human cognition-inspired Analogical Reasoning model performed the best. This success is likely due to explicitly computing state changes and comparing them with action options.
@@
Naive model Analogical Reasoning (AR)
@@
Figure 5: Probing where the model pays attention when connecting action & effects.. (a) Action is ‘putting something close to something’. Notice how the model tracks the box moved by the person to follow the state change. Note that the states and the action are from different samples. (b) Where the AR model attends vs where Naive model attends. AR model focuses on state changes and driving action and effects; notice the avocado and coin being tracked. Naive model without reasoning module seems to be doing simple matching based on texture. Zoom in to view better.
Human baseline. To benchmark our models, we assessed human performance. We had three undergraduate students answer a subset of 100 questions. Their average performance, detailed in Table 1, significantly exceeded that of all baseline models on this task of Action Selection.
Qualitative results. We analyze where our top-performing Analogical Reasoning model focuses when connecting actions & effects by extending the GradCAM method [41] (details on our visualization technique are in Appendix). The visualizations of the model’s attention during the Action Selection task are shown in Figure 5.
6.2 Learning visual representations from unlabeled videos using CATE
@@
Table 2: Nearest Neighbor retrieval quantitative results on UCF101. Best results: bolded; second best: underlined.
After pretraining with the self-supervised action selection task (section 4) on the UCF101 [42] train set without using any labels, we freeze the backbone (C3D architecture [47]). No fine-tuning is performed. The frozen backbone is then used as a feature extractor. Following standard practice in the literature [18, 49], we conduct a nearest neighbor retrieval experiment. Quantitative results for the nearest neighbor retrieval experiment are presented in Table 2. Our CATE model is compared with other state-of-the-art SSL video representation learning approaches. We observe that SSL-CATE performs the best overall in terms of Top-1 accuracy, ranking second best in the remaining metrics, only outperformed by STS. These results highlight the significance of CATE in effectively supporting the learning of general action representations.
We further probe into what the model might be learning by retrieving videos similar to the randomly chosen query videos. Results are presented in Figure 6. We observe that the retrieved results are mostly from the correct action class, suggesting that CATE encourages the model to learn the tracking of important human body parts and objects. Upon further investigation of some cases where the retrieval is from the incorrect action class, we found that motion patterns in the retrieved videos can potentially result in the same state change as the action in the query videos. For example, in archery and playing cello, pulling the arrow in archery and moving the bow/stick on the cello involve similar hand movements from humans. We further compare the results with a fully-supervised model. Qualitatively, we found the CATE-pretrained results to be on par with those from a fully supervised model.
6.3 SSL Effect-Affinity Assessment for AQA
In this case, we train using self-supervision grounded in connecting actions and their effects, as presented in section 5. Following the literature [40, 34], we carry out SSL training on the training set of the MTL-AQA dataset [36]. After SSL training, we do not perform any fine-tuning; we simply use the backbone as a feature extractor. Features are extracted from 16 uniformly sampled frames, and then a linear regressor is applied on top of them. We compare SSL CATE-Effect-Affinity Assessment with other self-supervised models, and the results are shown in Table 3. We observe that SSL-CATE Effect-Affinity Assessment self-supervision outperforms other state-of-the-art methods, indicating that CATE-Effect-Affinity Assessment does help in learning features suitable for fine-grained action analysis tasks such as action quality assessment.